### v1

# AdaSwitch: Adaptive Switching Generation for Knowledge Distillation

Abstract
Small language models (SLMs) are crucial for applications with strict latency and computational constraints, yet achieving high performance remains challenging. Knowledge distillation (KD) can transfer capabilities from large teacher models, but existing methods involve trade-offs: off-policy distillation provides high-quality supervision but introduces a training–inference mismatch, while on-policy approaches maintain consistency but rely on low-quality student outputs. To address these issues, we propose AdaSwitch, a novel approach that dynamically combines on-policy and off-policy generation at the token level. AdaSwitch allows the student to first explore its own predictions and then selectively integrate teacher guidance based on real-time quality assessment. This approach simultaneously preserves consistency and maintains supervision quality. Experiments on three datasets with two teacher–student LLM pairs demonstrate that AdaSwitch consistently improves accuracy, offering a practical and effective method for distilling SLMs with acceptable additional overhead.
AdaSwitch: Adaptive Switching Generation for Knowledge Distillation
Jingyu Peng‡§†, Maolin Wang§, Hengyi Cai, Yuchen Li, Kai Zhang11footnotemark: 1, Shuaiqiang Wang, Dawei Yin, Xiangyu Zhao§11footnotemark: 1 University of Science and Technology of China, § City University of Hong Kong, Baidu Inc. jypeng28@mail.ustc.edu.cn, morin.wang@my.cityu.edu.hk
1 Introduction
With the rapid development of large language models (LLMs), scaling laws suggest that increasing the number of parameters generally leads to better performance. However, small language models (SLMs) remain crucial in scenarios where latency and computational cost are highly sensitive. For instance, in search engines and recommender systems, it is common to process a large volume of queries simultaneously Li et al. (2025); Deng et al. (2025). Although LLMs may achieve higher accuracy, their high latency and computational demands can be prohibitive in such real-world applications. Moreover, SLMs also play an important role in accelerating inference methods such as speculative decoding Leviathan et al. (2023); Xia et al. (2024); Yin et al. (2024). Therefore, developing an SLM that achieves strong performance on a specific task is of vital importance.
Figure 1: An example illustrating the sequence generation process for off-policy, on-policy, and SKD methods. Refer to caption denotes the teacher, and Refer to caption represents the student.
Obtaining high-quality SLMs is often achieved by transferring knowledge from large models. Among various knowledge transfer techniques, knowledge distillation (KD) has emerged as a particularly effective approach Xu et al. (2024); Wang et al. (2025); Sanh et al. (2019). Unlike other transfer methods that may require extensive retraining on annotated datasets or rely solely on task-specific supervision, KD leverages the outputs of a teacher model to guide the student model Hinton et al. (2015). By utilizing the teacher’s predictions, including intermediate representations or logits, KD provides richer soft labels that capture nuanced information beyond hard ground-truth labels Yu et al. (2024); Aguilar et al. (2020). This enables the student to better approximate the teacher’s behavior, often leading to improved generalization, robustness, and efficiency, especially in scenarios with limited annotated data.
In KD, the selection of target sequences for sampling and supervision plays a pivotal role, as it directly shapes the quality and distribution of the learning signal. Based on this, KD methods are often categorized as off-policy or on-policy depending on the source of training data Xu et al. (2024). Off-policy approaches update the student model using either ground-truth sequences as hard labels or the teacher’s logits over sequences generated by the teacher Kim and Rush (2016). While this strategy ensures relatively high-quality training sequences, it can lead the student to suffer from a training–inference mismatch, as inference relies solely on the student’s own outputs. To address this issue, on-policy methods use the teacher’s logits on sequences generated by the student, thereby aligning the training process more closely with inference Agarwal et al. . However, due to the student’s limited capabilities, the quality of these student-generated sequences can be poor, especially in the early stages of distillation, restricting the effectiveness of distillation Xu et al. . Consequently, effectively combining the strengths of on-policy and off-policy approaches is key to unlocking the full potential of KD.
Most existing work focuses on data-level combinations, where sequences generated by the student and the teacher are simply mixed Lin et al. (2020); Agarwal et al. . However, this coarse-grained strategy cannot guarantee both high quality and consistency for each individual sequence. At the token-level, the only notable attempt is SKD, which introduces speculative decoding into sequence generation: the teacher supervises the student’s output at each token step Xu et al. . While this can align generation more closely with the teacher, the frequent switching between models produces “twisted” sequences that overfit to the teacher, yielding only marginal performance gains.
To address these limitations, we propose AdaSwitch, a two-stage sequence generation framework that dynamically alternates between on-policy and off-policy learning at the token level. In the exploration stage, the student generates an initial sequence autonomously, maintaining training–inference consistency while exploring with its latent capabilities. In the guidance stage, the teacher takes over and generates the remaining sequence to ensure high-quality outputs. Furthermore, the switching threshold is dynamically adjusted during generation, allowing the switch to adapt based on the difficulty of each sample. This one-time switch mechanism prevents excessive interventions, reduces sequence distortion and teacher overfitting, and maintains a balance between consistency and quality.
[…]
Figure 2: Overview of the proposed AdaSwitch approach. When the divergence between the student and teacher logits for the next token exceeds $\tau=K\cdot\bar{d}_{i}$, AdaSwitch switches to the teacher to generate the remaining tokens. Refer to caption denotes the teacher, and Refer to caption represents the student.
2 Methodology
2.1 Problem Setup
We study KD for LLMs by considering a teacher model $M_{t}$ and a student model $M_{s}$ with different capacities. Our goal is to train $M_{s}$ to mimic $M_{t}$’s behavior on a specific task $T$. The student model $M_{s}$ is parameterized by learnable weights $\theta_{s}$, while the teacher model $M_{t}$ has fixed parameters $\theta_{t}$.
For task $T$, we assume access to a collection of input prompts $x$ and associated target sequences $y$, either from ground-truth data or model-generated outputs. To quantify the difference between the teacher’s and student’s token-level predictions, we define a divergence function $D(\cdot\,\|\,\cdot)$, such as Kullback–Leibler (KL) divergence:
$d_{i}(M_{t},M_{s})=D\big(M_{t}(\cdot\mid y_{<i},x)\,\|\,M_{s}(\cdot\mid y_{<i},x)\big),$ (1)
where $i$ indexes the decoding step. The overall sequence-level divergence is then:
$D(M_{t}\|M_{s})(y|x)=\frac{1}{L_{y}}\sum_{i=1}^{L_{y}}d_{i}(x,y),$ (2)
with $L_{y}$ the length of $y$. The learning objective is to minimize $D(M_{t}\|M_{s})(y|x)$, encouraging $M_{s}$ to closely approximate $M_{t}$.
2.2 Target Sequence Selection
The choice of target sequence $y$ is critical in KD. Off-policy methods use fixed targets from datasets, while on-policy methods generate $y$ from the student model itself:
$y\sim M_{s}(\cdot\mid x;\theta_{s}).$ (3)
However, off-policy methods suffer from a mismatch between training and inference in the student model, whereas on-policy methods are limited by the low quality of student-generated sequences, leading to suboptimal performance.
Algorithm 1 AdaSwitch Knowledge Distillation for LLMs
1:Student LLM $M_{s}$, Teacher LLM $M_{t}$, Prompt dataset $\{x_{n}\}_{n=1}^{N}$, Decoding length $\alpha$, Sliding window length $L$, Divergence threshold multiplier $K$, Divergence metric $D$
2:for $n=1$ to NN do
3:  Initialize empty list of recent divergences $D_{\text{window}}$
4:  $use\_teacher\leftarrow\text{False}$
5:  for $i=1$ to $\alpha$ do
6:   if use_teacheruse\_teacher then
7:     Sample $y_{i}\sim M_{t}(\cdot\mid y_{<i},x_{n})$ $\triangleright$ teacher continues generating
8:   else
9:     Sample $y_{i}\sim M_{s}(\cdot\mid y_{<i},x_{n})$ $\triangleright$ student generates token
10:     Compute divergence $d_{i}=D\big(M_{t}(\cdot\mid y_{<i},x_{n})\,\|\,M_{s}(\cdot\mid y_{<i},x_{n})\big)$
11:     if $i>L$ then
12:      Compute sliding window average $\bar{d}_{i-1}=\frac{1}{L}\sum_{j=i-L}^{i-1}d_{j}$
13:      if $d_{i}>K\cdot\bar{d}_{i-1}$ then
14:        $use\_teacher\leftarrow\text{True}$ $\triangleright$ switch to teacher for remaining tokens
15:        Sample $y_{i}\sim M_{t}(\cdot\mid y_{<i},x_{n})$
16:      end if
17:     end if
18:     Append $d_{i}$ to $D_{\text{window}}$
19:   end if
20:  end for
21:  Apply gradient descent to minimize $D(M_{t}\|M_{s})(y\mid x_{n})$ over generated sequence
22:end for
2.3 AdaSwitch
To overcome the limitations of off-policy and on-policy approaches and fully unlock the potential of knowledge distillation, we propose a novel sequence generation approach known as AdaSwitch as shown in Figure 2.
Adaswitch is designed to enhance the training of student models during sequence generation through dynamic guidance based on the divergence between the predictions made by the student and the teacher models. The primary objective of AdaSwitch is to achieve a delicate balance between exploration and guidance: while the student model is permitted to explore early predictions, it receives corrective interventions when its outputs diverge significantly from those of the teacher, thereby ensuring high-quality supervision throughout the training process.
At each generation step $i$, we compute the divergence $d_{i}(M_{t},M_{s})$ between the next-token predictions from the teacher and the student, as defined in Equation 1. This divergence serves as a quantitative measure of the discrepancy between the two models’ predictions on the current token, allowing us to evaluate the alignment of the student with the teacher’s knowledge.
To effectively assess whether the current predictions of the student model diverge from those of the teacher model, we introduce a metric that adapts to sequences characterized by varying distributions. Specifically, we maintain a sliding window of length $L$ over the most recent divergences, computing the moving average as follows:
$\bar{d_{i}}=\frac{1}{L}\sum_{j=i-L+1}^{i}d_{j},\quad\text{for }i\geq L.$ (4)
This moving average, $\bar{d_{i}}$, provides a context-sensitive measure of divergence, allowing the model to adjust its behavior according to recent prediction patterns. Based on this, we define a threshold $\tau$ when generating $y_{i}$ as follows:
$\tau=K\cdot\bar{d}_{i-1},$ (5)
where $K$ is a hyperparameter that scales the threshold relative to the moving average of divergences. The student model continues generating tokens as long as its divergence from the teacher model remains below this adaptive threshold $\tau$. If the divergence exceeds $\tau$, the teacher continues to generate:
$y_{i}\sim\begin{cases}M_{s}(\cdot\mid y_{<i},x;\theta_{s}),&d_{i}(M_{t},M_{s})\leq\tau,\\ M_{t}(\cdot\mid y_{<i},x;\theta_{t}),&d_{i}(M_{t},M_{s})>\tau.\end{cases}$ (6)
Once a switch to the teacher model occurs, all subsequent tokens are generated exclusively by the teacher model, eliminating the need for further divergence checks. The complete algorithm of AdaSwitch is presented in Algorithm 1.
By integrating adaptive sliding-window thresholds and selective teacher intervention, AdaSwitch not only ensures consistency between training and inference but also maintains high-quality supervision for the student model. This approach effectively harnesses the strengths of both exploration and guidance, leading to improved model performance and robustness in diverse sequence generation tasks.
Llama-3.1-3b to Llama-3.1-1b Qwen-2.5-3b to Qwen-2.5-0.5b
SUMM GSM GSM_Plus SUMM GSM GSM_Plus
Teacher 37.45 59.59 38.58 37.41 65.35 45.42
Student 13.48 0.00 0.00 12.07 0.00 0.00
SFT 32.34 29.04 12.00 30.28 31.01 15.54
SeqKD 32.09 28.20 11.78 30.13 30.42 14.42
Off-policy KD 34.57 30.55 13.83 32.44 33.51 17.92
On-policy KD 34.34 32.68 14.33 32.31 34.12 20.17
ImitKD 35.07 30.55 15.63 32.50 34.57 21.29
SKD 34.88 32.37 16.88 32.27 32.98 20.38
AdaSwitch 35.56 35.03 17.33 32.60 36.32 21.63
Table 1: Performance comparison of different Knowledge Distillation methods for Llama and Qwen on three tasks. Performance is evaluated using Rouge-L on the SUMM summarization dataset and accuracy on the GSM and GSM-Plus reasoning benchmarks.
3 Experiments
3.1 Student and Teacher Models
We conducted experiments using two state-of-the-art LLM families: Qwen 2.5 Team (2024) and Llama 3.1 Grattafiori et al. (2024). For the teacher models, we employed the SFTed Qwen-2.5-3B and Llama-3.1-3B, while the student models consisted of Qwen-2.5-0.5B and Llama-3.1-1B.
3.2 Datasets and Experimental Setup
To evaluate the effectiveness of our proposed AdaSwitch, we conduct experiments on two distinct types of datasets: DialogSum Chen et al. (2021) for dialogue summarization and GSM Cobbe et al. (2021) for arithmetic reasoning. For arithmetic reasoning, we use one additional dataset GSM-Plus Li et al. (2024) to further test the generalizability and robustness of our approach. Each task follows a two-stage procedure: we first fine-tune the teacher model on the task, under the assumption that it serves as a domain expert, and then perform knowledge distillation by randomly sampling approximately 1K examples from the task, following common practices in prior KD work. Detailed configurations for each dataset are provided in the subsequent sections. For reproducibility, all implementations adopt greedy decoding during evaluation. Detailed descriptions of the datasets, including metrics and data splits, are provided in Appendix A.
3.3 Hyperparameters
For all SFT processes, we used a learning rate of 1e-5, a warmup ratio of 0.1, and a dropout rate of 0.1. Each SFT checkpoint was trained for three epochs, with the checkpoint exhibiting the lowest validation loss being selected. For the KD process, we set the maximum input and output lengths to {1024, 256} for dialogue summarization and {128, 512} for arithmetic reasoning. The student models were configured with a temperature of 0.5 and a top-p of 0.5, while the teacher models had a temperature of 0.2 and a top-p of 0.5, as per previous research. All tasks and models were trained for three epochs with a learning rate of 1e-5. To ensure a fair comparison, the $K_{skd}$ for SKD was set to 0.5, as reported in its original paper. We set the sliding window length $L$ to 10 and the threshold $K$ to 3 without specific fine-tuning. The impact of these hyperparameters is discussed later.
3.4 Baselines
Supervised FT directly minimize the negative log-likelihood of the student’s predictions on the target sequence, which is the ground-truth.
SeqKD Kim and Rush (2016) performs Supervised FT on teacher-generated target sequences.
SupervisedKD Sanh et al. (2019) trains a student model to mimic the token-level probability distribution of a teacher model over a fixed ground-truth sequence of each $(x,y)$ pair by minimizing the distance between the student’s and teacher’s predictions. Since this method delivers the best performance among all off-policy approaches, we refer to it simply as off-policy in the following.
GKD Agarwal et al. applied on-policy KD to address the training-inference mismatch in student models by sampling target tokens directly from the student’s own output. We refer to this method as on-policy in the remainder of this paper.
ImitKD Lin et al. (2020) randomly sample target sequences from either the ground truth dataset or the student model, which is a sequence-level mixing strategy of on-policy and off-policy.
SKD Xu et al. employs a speculative decoding method to improve the quality of the generated sequences by using the teacher model to supervise the sequence generation.
Figure 3: Further comparison of performance under different distance metrics on three tasks.
3.5 Overall Performance
As shown in Table 1, on-policy and off-policy approaches exhibit complementary strengths. On-policy distillation outperforms on GSM and GSM-Plus, whereas off-policy achieves superior results on SUMM. These observations suggest that integrating both methods can be advantageous, allowing the student model to benefit from the strengths of each approach.
Mixed methods such as ImitKD and SKD attempt to integrate both paradigms and yield localized improvements. For instance, ImitKD performs well on SUMM, and SKD shows gains on GSM-Plus. However, these improvements are marginal and inconsistent across tasks and model pairs, reflecting the limited performance improvement of their designs.
In contrast, AdaSwitch achieves state-of-the-art results across all benchmarks, demonstrating both robustness and effectiveness. On GSM, AdaSwitch surpasses the second-best method by 7.2% and 11.8% when applied to Llama and Qwen, respectively. These results highlight its ability to integrate the strengths of on-policy and off-policy learning at the token level, leading to significant and consistent improvements.
For a more comprehensive comparison, we evaluate performance using JSD and reversed KLD Gu et al. (2024) as distance metrics, as shown in Figure 3. The results indicate that, while existing mixed KD methods occasionally provide improvements, their performance is inconsistent. By comparison, although results vary slightly across different metrics, AdaSwitch consistently performs well and is compatible with different commonly used distance measures, highlighting the robustness and flexibility of AdaSwitch in diverse scenarios.
3.6 Analysis on Switch Rate and KLD
To better understand the distillation process of AdaSwitch, we analyze the switch rate, defined as the proportion of sequences that trigger a switch from exploration to guidance, along with the average KLD between the student and teacher at the switching points. Statistics were collected every 100 steps, and the results are shown in Figure 4.
First, we observe a significant decrease in KLD during distillation across all tasks. This trend indicates that the student model gradually converges toward the teacher, consistent with intuitive expectations. Moreover, the KLD in the SUMM task is consistently lower than that in GSM, aligning with the performance difference between the two tasks and supporting the view that SUMM is relatively easier than GSM.
Switch rates also reveal task-dependent differences. On GSM, switch rates remain above 95% throughout, whereas SUMM exhibits substantially lower rates. This suggests that more challenging tasks demand stronger reliance on teacher guidance, while easier tasks allow greater exploration. AdaSwitch thus adaptively balances exploration and guidance according to task difficulty.
Further, we observe that the switch rate increases steadily during distillation, most notably in SUMM, while remaining near 100% in GSM. This trend reflects a gradual decline in exploration and a rise in guidance, indicating that on-policy distillation contributes less over time. Although this may seem counterintuitive, since weaker students would normally require more guidance, the pattern is explained by KLD: early in training, high divergence suppresses switching, but as the student aligns with the teacher, switching becomes more frequent. In other words, AdaSwitch begins in an exploration-dominated on-policy mode and gradually shifts to off-policy guidance as the student improves. This adaptive behavior allows AdaSwitch to dynamically balance exploration and guidance according to the student’s capability, ensuring optimal distillation across varying tasks. Due to space limitations, we provide the complete experimental results and the analysis of relative switch positions during distillation in Appendix B.
3.7 Early-Stage Distillation Analysis
The challenge of low-quality outputs in on-policy KD is particularly acute during the early stages of the distillation process. To investigate whether AdaSwitch addresses this issue, we present the validation loss and test performance trends for on-policy, off-policy, and AdaSwitch methods over the first 100 steps in Figure 5.
Regarding the validation loss trends, the result indicates that AdaSwitch consistently achieves values intermediate between those of the on-policy and off-policy methods, providing evidence of an effective integration. Since validation loss is computed using the teacher model’s outputs as ground truth, the off-policy method, which directly optimizes against these outputs, attains the lowest validation loss.
Further experimental results of the performance on the test set reveals that AdaSwitch demonstrates the most rapid improvement rate, particularly on the GSM dataset. Notably, at step 30, the accuracy of AdaSwitch surpasses 10%, whereas the off-policy and on-policy methods achieve only 7.96% and 0%, respectively. These results suggest that AdaSwitch, through effective token-level mixing, simultaneously ensures the quality of generated sequences and reduces the discrepancy between training and inference sequences during the early stages of distillation. The full experimental results are presented in Appendix B.
Figure 4: Analysis of the switch rate and the KLD at the switch token between the student and teacher models throughout the distillation process. Statistics were collected every 100 steps for the first 1000 steps.
3.8 Parameter Analysis
In the proposed AdaSwitch framework, two hyperparameters, $K$ and $L$, play pivotal roles: $K$ regulates the threshold for switching between exploration and guidance modes, while $L$ determines the window size for this switching mechanism.
Regarding $K$, empirical results indicate that optimal performance is achieved when $K$ lies between 3 and 4. Increasing $K$ tends to elevate the proportion of exploration relative to guidance, whereas decreasing $K$ has the opposite effect. When comparing performance across datasets, we observe that the impact of varying $K$ is relatively minor and more stable on the SUMM dataset. In contrast, on the GSM dataset, particularly for the Llama model, performance is significantly more sensitive to changes in $K$.
For the parameter $L$, the best results are generally obtained around a value of 10. Larger values of $L$ increase the minimum duration of exploration (i.e., at least $L$ tokens), while simultaneously reducing the sensitivity of the switching mechanism to fluctuations in the KLD. Conversely, smaller $L$ values increase sensitivity to KLD changes. Consistent with observations for $K$, the effects of varying $L$ are more stable and less pronounced on the SUMM dataset, whereas on the GSM dataset–and especially with the Llama model–performance exhibits greater susceptibility to $L$ variations.
Figure 5: Performance on the validation and test sets during the early stages of distillation.
3.9 Time Consumption Analysis
In AdaSwitch, the time cost primarily depends on the sequence length and the per-token generation speed. However, in practice, it is difficult to estimate the sequence length for each sample across different KD methods, since generation is performed by different models. Moreover, sequence generation accounts for only a small fraction of KD time. Therefore, we conducted a general and practical experiment by comparing the total runtime of different KD methods over the distillation.
The results in Table 2 indicate that off-policy consistently achieves the lowest runtime cost, since it doesn’t conduct sequence generation. On the other hand, both SKD and AdaSwitch are slower than on-policy because they require computing teacher logits during generation and switching between the student and teacher.
[…]
Figure 6: Parameter analysis on $K$ and $L$.
4 Related Work
4.1 Small Language Model Optimization
The scaling law suggests that increasing the number of parameters makes a language model more powerful. However, relatively small language models (SLMs) are essential in scenarios that are sensitive to latency or computational costs, such as search engines Li et al. (2025) and recommendation systems Deldjoo et al. (2024). Therefore, developing a small language model with satisfactory performance for specific tasks becomes crucial.
Various optimization methods can be employed to create an SLM, including Supervised Fine-Tuning (SFT), Direct Policy Optimization (DPO) Rafailov et al. (2023), and Knowledge Distillation (KD) Hinton et al. (2015). However, both SFT and DPO require large amounts of high-quality labeled data Raghavendra et al. (2025); Wang et al. (2024), which can be a significant limitation. In contrast, KD can leverage data of arbitrary size by utilizing a teacher model, making it more scalable. Moreover, while SFT and DPO generally rely on hard labels, KD has the advantage of using intermediate representations or logits from the teacher model as supervisory signals, providing richer information for training the student model Kim et al. (2022). Given the advantages and widespread use of KD for training SLMs Muralidharan et al. (2024); Guo et al. (2025); Yang et al. (2025); Abdin et al. (2024), consequently, our work builds on KD to refine its core mechanism for improved performance.
4.2 Knowledge Distillation for LLMs
In the era of LLM, KD has been adapted to autoregressive models and categorized into off-policy and on-policy methods, inspired by reinforcement learning Xu et al. (2024). Off-policy KD uses either ground-truth sequences or teacher-generated logits for supervision Kim and Rush (2016); Sanh et al. (2019). While this ensures high-quality training data, it often causes a mismatch between the student’s training distribution and inference-time behavior, leading to inconsistencies Agarwal et al. . Conversely, on-policy KD trains the student using its own generated sequences, improving consistency and performance on some tasks. However, the student’s initially poor outputs limit sequence quality and thus KD effectiveness Xu et al. . Prior attempts to combine off- and on-policy KD typically mix data at a coarse level Lin et al. (2020); Agarwal et al. . The most related approach, SKD, employs speculative decoding guided by teacher supervision to enhance sequence quality Xu et al. . However, its frequent switching between student and teacher models during generation can cause overfitting to the teacher and restrict performance improvements. In contrast, AdaSwitch mitigates this issue by introducing a more stable and efficient token-level mixing strategy.
SUMM GSM
Llama Qwen Llama Qwen
Off-policy 0.77x 0.79x 0.61x 0.57x
On-policy 1.00x 1.00x 1.00x 1.00x
SKD 1.28x 1.31x 1.61x 1.53x
AdaSwitch 1.20x 1.22x 1.41x 1.38x
Table 2: Relative runtime cost of different KD methods over the distillation process. Results are normalized by On-policy (set as 1.00x)
5 Conclusion
In this paper, we introduce AdaSwitch, a token-level mixed distillation framework designed to reduce the gap between training and inference while preserving high-quality student outputs. AdaSwitch adopts a two-stage generation process that dynamically switches from student exploration to teacher guidance. Extensive experiments on three tasks, two LLM families, and three distance metrics verify its effectiveness, delivering significant performance gains with acceptable time cost.
6 Limitations
Our primary limitation lies in the constrained scope of our experimental evaluation. We validate AdaSwitch on three tasks and two LLM families, which may not fully capture the breadth of challenges faced in diverse real-world applications. Future work will extend the evaluation to a wider range of tasks and domains to further verify the generality and robustness of the proposed method. Additionally, while AdaSwitch achieves significant performance gains, it introduces an additional computational overhead during the KD process. As shown in our time complexity analysis, AdaSwitch incurs approximately 20% higher time consumption compared to the standard on-policy KD method. However, we consider this trade-off acceptable because the KD process is typically performed offline, and crucially, our method introduces no additional latency or computational cost during the inference phase for the final student model.
Appendix A Datasets
Dialogue Summarization. Chen et al. (2021) For the dialogue summarization task, we employ the DialogSum dataset, which contains 12.5K training samples, 500 validation samples, and 1.5K test samples. The complete training set is utilized for SFT of the teacher models. For KD, we randomly sample 1,000 instances from the training data, and the effectiveness of the KD methods is subsequently evaluated on the held-out test set. To assess summary quality, we use the ROUGE-L metric.
GSM8K. Cobbe et al. (2021) To evaluate arithmetic reasoning, we utilized the GSM8K dataset, a common benchmark for assessing how well language models solve math problems. The dataset comprises 8,500 training and 1,300 test examples, with each example presenting a grade-school word problem alongside its step-by-step solution and final answer. We partitioned the GSM8K training set into a training subset (7,000 instances) for SFT of the teacher models and a validation subset (473 instances) following  Xu et al. . For KD, we randomly selected 1,000 problems from the training data and then evaluated the effectiveness of our KD methods on the separate test set. Accuracy is applied to assess the performance.
GSM-Plus. Li et al. (2024) GSM-Plus is an adversarial dataset that evaluates the robustness of large language models on mathematical reasoning. This challenging benchmark modifies the original GSM8K problems with numerical, arithmetic, and textual perturbations to test if models can maintain correct reasoning despite subtle changes. In our experiments, we evaluated the student models, which were distilled on GSM8K, on the GSM-Plus test set to further assess the generalizability and robustness of our knowledge distillation methods. The model’s performance is assessed using accuracy.
Figure 8: Analysis of the validation loss in early stage of distillation
Appendix B More Experimental Results
We provide the complete experimental results for the Analysis of Switch Rate and KLD, as well as the Early-Stage Distillation Analysis, in Figure 7 and Figure 8, respectively.
Moreover, we analyze the relative switching position of AdaSwitch during distillation. This also reflects the proportion of tokens generated by the student model. As shown in Figure 9, in GSM the switching position exhibits a clear upward trend, indicating that the teacher intervenes later in the generation process. This suggests that the student model is gradually approaching the teacher’s capability. In contrast, for the SUMM task we do not observe a noticeable increase, which we attribute to the smaller performance gap between the student and teacher models in this setting.
Figure 9: Analysis of the relative switching position of distillation process